OpenStudy (anonymous):

HELP: Simplify: (3-2i)(4+3i)

OpenStudy (matt101):

(3-2i)(4+3i) = 12 + 9i - 8i - 6i^2 = 12 + i + 6 = 18 + i *Remember that i^2 = -1

OpenStudy (anonymous):

so what's the answer then

OpenStudy (anonymous):

im a bit confused on this

OpenStudy (matt101):

The answer is 18 + i. Your can't simplify it any further because i is an imaginary number equal to the square root of -1.

OpenStudy (anonymous):

oh okay you sure? LOL

geerky42 (geerky42):

(3-2i)(4+3i) Do FOIL. You know that, right?

geerky42 (geerky42):

Or you can do box method.

OpenStudy (anonymous):

yes, first, outside, inside, last okay thanks males